# Liouville polarization and the rigidity of their Lagrangian skeleta in dimension $4$

### v1

@@ 4.1.2. Reduction to the monotone setting
be the extendable Liouville polarization of $D(A)\times D(B)$ with skeleton $\Gamma_{1}\times\Gamma_{2}$ provided by Lemma 3.5. We recall that in this case, an extension is obtained by taking somewhat larger discs, so $\overline{\Sigma}$ is just the closure of $\Sigma$ in $\mathbb{C}^{2}$, i.e. the closed discs. Similarly, let ${\mathcal{P}}^{\prime}:=\{p_{1}^{\prime},\dots,p_{m}^{\prime}\}\subset D(ma)$, ${\mathcal{Q}}^{\prime}:=\{q_{1}^{\prime},\dots,q_{n}^{\prime}\}\subset D(nb)$, and let
@@
be the Liouville polarization of $D(ma)\times D(nb)$ with skeleton $\Gamma^{\prime}_{\!1}\times\Gamma^{\prime}_{\!2}$. Since we are free to choose $\Gamma^{\prime}_{\!1}$ and $\Gamma^{\prime}_{\!2}$, and since $D(A)\subset D(ma)$ and $D(B)\subset D(nb)$, we may as well assume that $p_{i}^{\prime}=p_{i}$ and $q_{j}^{\prime}=q_{j}$. This is not an important assumption but will simplify the proof at some point. We will nevertheless keep the primes in order to distinguish the source and the target. Lemma 3.6 provides disjoint open discs $D_{i}^{A}$ in $D(ma)\setminus\Gamma_{1}$, disjoint open discs $D_{j}^{B}$ in $D(nb)\setminus\Gamma_{2}$, points $p_{i}^{\prime 1},p_{i}^{\prime 2}\in D_{i}^{A}$ for $i\leq m^{\prime}$, $q_{j}^{\prime 1},q_{j}^{\prime 2}\in D_{j}^{B}$ for $j\leq n^{\prime}$, and a Liouville polarization $({\bf\Sigma}^{\prime\prime},\lambda^{\prime\prime})$ of $D(ma)\times D(nb)$ with
[…]
@@
${\mathcal{B}}\left(\cup_{i}D_{i}^{A}\times D(nb)\bigcup\cup_{j}D(ma)\times D_{j}^{B},\lambda^{\prime\prime}\right)$
@@
Case $B<nb$: As already noticed, in this case $\sigma$ can be extended to an area preserving embedding $\hat{\sigma}\colon D(B+\varepsilon)\to D(nb)$, so $\phi$ can be extended to a smooth area preserving embedding $\widehat{\Sigma}\to\Sigma^{\prime\prime}$ which is a symplectic morphism between our polarizations. It is $(\alpha_{\operatorname{st}},\alpha_{\operatorname{st}})$-exact because all components of $\Sigma$ are discs. Theorem 2.20 now implies that
@@
${\bf\Sigma}^{\prime\prime}\,\subset\,\cup_{i}D_{i}^{A}\times D(nb)\bigcup D(ma)\times\cup_{j}D_{j}^{B},$
@@
${\mathcal{B}}({\bf\Sigma}^{\prime\prime},\lambda^{\prime\prime})$$\subset$ ${\mathcal{B}}\Bigl{(}\cup_{i}D_{i}^{A}\times D(nb)\bigcup D(ma)\times\cup_{j}D_{j}^{B}\Bigr{)}$
@@
Composing (4.1.2) with (4.1.2) we obtain the asserted exact symplectic embedding.
Case $B=nb$: In this case, $\sigma=\mbox{id}$ does not extend to a larger disc with image still in $D(nb)$, so formally the above argument does not apply. We solve this difficulty by revisiting the proof of Theorem 2.20 in this particular case: Although $\phi$ does not extend to an extended polarization $\widehat{\Sigma}$ as before, the product map $\psi:=\tau\times\mbox{id}$ extends $\phi$ to a small neighbourhood of $\Sigma$ in $D(A)\times D(b)$, still with image in $D(ma)\times D(nb)$. The forms $\lambda$ on $D(A)\times D(B)$ and $\lambda^{\prime\prime}$ on $D(ma)\times D(nb)$ are both split:
@@
where $\lambda_{1},\lambda_{2}$ are defined on $D(A)\backslash\{p_{i}\}$, $D(B)\backslash\{q_{j}\}$ and $\lambda_{1}^{\prime\prime},\lambda_{2}^{\prime\prime}$ are defined on $D(ma)\backslash\{p_{i}^{\prime},p_{i}^{\prime 2}\}$, $D(mb)\backslash\{q_{j}^{\prime}\}$. Since $B=mb$ and $q_{j}=q_{j}^{\prime}$ we can take $\lambda_{2}=\lambda_{2}^{\prime\prime}$, and since $\lambda_{1}$ and $\lambda_{1}^{\prime\prime}$ have the same residues at $p_{i}$ and $p_{i}^{\prime}=p_{i}$, we can make them coincide on a small neighbourhood of the $p_{i}$. Since $\tau$ can clearly be taken to be the identity on a neighbourhood of the points $p_{i}$, the map $\psi=\tau\times\mbox{id}$ (defined on a neighbourhood of $\Sigma$) then pulls back $\lambda^{\prime\prime}$ to $\lambda$. Thus the basic conjugacy procedure described in the proof of Theorem 2.20 to embed ${\mathcal{B}}({\bf\Sigma},\lambda)$ into ${\mathcal{B}}({\bf\Sigma}^{\prime\prime},\lambda^{\prime\prime})$ applies, and now the argument in the previous case applies. $\square$
@@ 4.2. Proof of Theorem 2
where $\lambda_{z}=\lambda_{w}$ is a Liouville form on $\mathbb{R}^{2}\backslash\bigl{(}\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})\bigr{)}$, with $\Gamma_{1}$ as skeleton. We construct this Liouville form exactly as in Section 3.1, where compactness plays no role. Alternatively, we can do the construction in the proof of Proposition 3.2 on $[-2,2]^{2}$ in a $\mathbb{Z}^{2}$-periodic way. We then obtain a smooth Liouville form on $[0,1]^{2}$ that extends to a smooth $\mathbb{Z}^{2}$-periodic Liouville form on $\mathbb{R}^{2}\backslash\bigl{(}\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})\bigr{)}$ with skeleton $\Gamma_{1}$. $\square$
@@
    Then $\bigl{(}(\Sigma_{N},1),\lambda\bigr{)}$ is an extendable Liouville polarization of $]-N,N[^{4}$ whose skeleton is $\Gamma_{N}$. Notice that $\Sigma_{N}\subset\Sigma_{N+1}$.
@@
    Then $\bigl{(}(\Sigma_{Z},1),\lambda\bigr{)}$ is a Liouville polarization of $Z(2)$.
Now $\Sigma_{Z}$ is made of two planes {(±$\{(\pm\frac{1}{2},\frac{1}{2})\}\times\mathbb{R}^{2}$ and infinitely many discs
@@
($\bigl{(}\{(\Sigma^{\prime}_{1},1),(\Sigma^{\prime}_{2},1)\},\lambda^{\prime}\bigr{)}$
is a Liouville polarization of $Z(2)$.
@@
For each $N$ there exists a symplectic morphism $\phi_{N}\colon\widehat{\Sigma}_{N}\to\Sigma^{\prime}$ such that
@@
  • $\Sigma_{2}^{\prime}$ is obtained by gluing $(\frac{1}{2},\frac{1}{2})\times\mathbb{R}^{2}$ with the topological discs $D_{n,m}$, $(n,m)\in\mathbb{Z}^{2}$. This can be done by replacing each round disc $(\frac{1}{2},\frac{1}{2})\times D\bigl{(}(n+\frac{1}{2},m+\frac{1}{2}),\varepsilon\bigr{)}$ by its gluing with the disc $D_{n,m}$. These gluings can be parameterized by $\mathbb{Z}^{2}$-translations of a punctured disc $D(0,\varepsilon)\backslash\{(0,\frac{\varepsilon}{2}\}$ with an area form $\omega_{\varepsilon}$ whose total area is $2+\varepsilon$, and that is standard near the boundary $S^{1}(\varepsilon)$. Each point $I_{n,m}$ is again parameterized by $(n+\frac{1}{2},m+\frac{1}{2})$.
We also define $\Sigma_{1}^{\prime}(N):=\{(-\frac{1}{2},\frac{1}{2})\}\times[-N,N]^{2}$ and $\Sigma_{2}^{\prime}(N)\subset\Sigma_{2}^{\prime}$ the part of $\Sigma_{2}^{\prime}$ parameterized by $[-N,N]^{2}\backslash(\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{2}))$, and notice that they correspond exactly to the intersections of these curves with $Z(2)\cap[-N,N]^{4}$ (see Figure 4.2).
@@
Figure 4.2. The parametrization of the curve $\Sigma_{2}^{\prime}$. Each depicted disc $D(\varepsilon)$ has area $\varepsilon$ on the picture, but carries the symplectic form $\omega_{\varepsilon}$ of area $2+\varepsilon$. The black points are the punctures, and the red crosses are the intersection points with $\Sigma_{1}^{\prime}$.
@@
We construct $\phi_{N}$ by sending a slightly larger disc around the closure of each $\Sigma_{N}^{V}(n,m)$ into $\Sigma_{1}^{\prime}(2N^{2}+1)$, and by sending a slightly larger disc around the closure of each $\Sigma_{N}^{H}(n,m)$ into $\Sigma^{\prime}_{2}(2N^{2}+1)$. The areas of these surfaces are
@@
#$\#\bigl{(}\Sigma_{1}^{\prime}(2N^{2}+1)\cap\Sigma_{2}^{\prime}(2N^{2}+1)\bigr{)}=(2(2N^{2}+1))^{2}>16N^{4}=\#(\Sigma_{N}^{V}\cap\Sigma_{N}^{H}).$
We can therefore define $\phi_{N}$ by its restrictions to $\Sigma_{1}^{V}$ and $\Sigma_{1}^{H}$ as illustrated in Figures 4.3 and 4.4, where $\phi_{N}$ is explained in blue and $\phi_{N+1}$ in yellow for $N=1$. These are area preserving embeddings that take the points at the red crosses to points at the red crosses. This defines symplectic morphisms $\phi_{N}\colon\widehat{\Sigma}_{N}\to\Sigma^{\prime}$, and we can clearly choose $\phi_{N+1}$ such that $\phi_{N+1}|_{\widehat{\Sigma}_{N}}=\phi_{N}$. $\square$
@@
Figure 4.3. The restriction of $\phi_{1}$ to $\Sigma_{1}^{V}$ and of $\phi_{2}$ to $\Sigma_{2}^{V}$. Each band has height slightly less than $1$, so that we can construct a larger band of height less than $1$ in the next step $N+1$.
@@
Figure 4.4. The restriction of $\phi_{1}$ to $\Sigma_{1}^{H}$ and of $\phi_{2}$ to $\Sigma_{2}^{H}$. These are almost the same embeddings as in the previous figure, except that they avoid the punctures. Note that the area $2+\varepsilon$ of the discs is not even used since for $\varepsilon>0$ small enough the area of $\Sigma_{2}^{\prime}(2N^{2}+1)$ without the discs is already larger than the area of $\Sigma_{N}^{H}$.
@@
defined on a neighourhood $\widehat{U}$ of $\Sigma$ such that $\Phi(\widehat{U})\subset\psi(U)$, whence the closure of $\Phi(\widehat{U})$ is also disjoint from $\partial Z(2)$. Choosing $\widehat{U}$ smaller if necessary, we can also assume that $\widehat{U}$ retracts onto $\Sigma$, whose components are topological discs. We therefore find a Liouville form $\lambda^{\prime\prime}$ on $Z(2)$ that coincides with $\Phi_{*}\lambda$ on $\Phi(\widehat{U})$ and with $\lambda^{\prime}$ near $\partial Z(2)$. Then the map
@@
Since the components of $\Sigma$ are discs, the $\phi_{N}$ are $(\alpha_{{\operatorname{st}}}|_{\Sigma_{N}},\alpha_{{\operatorname{st}}}|_{\Sigma^{\prime}})$-exact, and so $\Psi_{1}$ is $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact. The asserted $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact symplectic embedding $\Psi\colon{\mathcal{B}}(\Sigma,\lambda)=\mathbb{R}^{4}\setminus\Gamma\to Z(2)$ is now obtained from Lemma 2.21. $\square$

### v2

@@ 4.2.2. Reduction to the monotone setting
be the extendable Liouville polarisation of $D(A)\times D(B)$ with skeleton $\Gamma_{1}\times\Gamma_{2}$ provided by Lemma 3.5. We recall that in this case, an extension is obtained by taking somewhat larger discs, so $\overline{\Sigma}$ is just the closure of $\Sigma$ in $\mathbb{C}^{2}$, i.e. the closed discs. Similarly, let ${\mathcal{P}}^{\prime}:=\{p_{1}^{\prime},\dots,p_{m}^{\prime}\}\subset D(ma)$, ${\mathcal{Q}}^{\prime}:=\{q_{1}^{\prime},\dots,q_{n}^{\prime}\}\subset D(nb)$, and let
@@
be the Liouville polarisation of $D(ma)\times D(nb)$ with skeleton $\Gamma^{\prime}_{\!1}\times\Gamma^{\prime}_{\!2}$. Since we are free to choose $\Gamma^{\prime}_{\!1}$ and $\Gamma^{\prime}_{\!2}$, and since $D(A)\subset D(ma)$ and $D(B)\subset D(nb)$, we may as well assume that $p_{i}^{\prime}=p_{i}$ and $q_{j}^{\prime}=q_{j}$. This is not an important assumption but will simplify the proof at some point. We will nevertheless keep the primes in order to distinguish the source and the target. Lemma 3.6 provides disjoint open discs $D_{i}^{A}$ in $D(ma)\setminus\Gamma_{1}$, disjoint open discs $D_{j}^{B}$ in $D(nb)\setminus\Gamma_{2}$, points $p_{i}^{\prime 1},p_{i}^{\prime 2}\in D_{i}^{A}$ for $i\leq m^{\prime}$, $q_{j}^{\prime 1},q_{j}^{\prime 2}\in D_{j}^{B}$ for $j\leq n^{\prime}$, and a Liouville polarisation $({\bf\Sigma}^{\prime\prime},\lambda^{\prime\prime})$ of $D(ma)\times D(nb)$ with
${\bf\Sigma}^{\prime\prime}$ $\,:=\,\Bigl\{$ $\bigl(p_{i}^{\prime 1}\times D(nb),a_{i}\bigr)_{i\leq m^{\prime}},\bigl(p_{i}^{\prime 2}\times D(nb),a-a_{i}\bigr)_{i\leq m^{\prime}},$
@@
${\mathcal{B}}\Bigl(\bigcup_{i}D_{i}^{A}\times D(nb)\cup\bigcup_{j}D(ma)\times D_{j}^{B},\lambda^{\prime\prime}\Bigr)$
@@
Case $B<nb$: As already noticed, in this case $\sigma$ can be extended to an area preserving embedding $\hat{\sigma}\colon D(B+\varepsilon)\to D(nb)$, so $\phi$ can be extended to a smooth area preserving embedding $\widehat{\Sigma}\to\Sigma^{\prime\prime}$ which is a symplectic morphism between our polarisations. Its restriction to each component of $\Sigma$ is $(\alpha_{\operatorname{st}},\alpha_{\operatorname{st}})$-exact, since all components are discs. By Lemma 4.2, $\phi$ can be modified to an $(\alpha_{\operatorname{st}},\alpha_{\operatorname{st}})$-exact symplectic morphism, and Theorem 2.20 now implies that
@@
${\bf\Sigma}^{\prime\prime}\,\subset\,\bigcup_{i}D_{i}^{A}\times D(nb)\cup\bigcup_{j}D(ma)\times D_{j}^{B},$
@@
${\mathcal{B}}({\bf\Sigma}^{\prime\prime},\lambda^{\prime\prime})$ $\subset$ ${\mathcal{B}}\Bigl(\bigcup_{i}D_{i}^{A}\times D(nb)\cup\bigcup_{j}D(ma)\times D_{j}^{B}\Bigr)$
@@
Composing (4.2) with (4.2.2) we obtain the asserted exact symplectic embedding.
Case $B=nb$: In this case, $\sigma=\mbox{id}$ does not extend to a larger disc with image still in $D(nb)$, so formally the above argument does not apply. We solve this difficulty by revisiting the proof of Theorem 2.20 in this particular case: Although $\phi$ does not extend to an extended polarisation $\widehat{\Sigma}$ as before, the product map $\psi:=\tau\times\mbox{id}$ extends $\phi$ to a small neighbourhood of $\Sigma$ in $D(A)\times D(B)$, still with image in $D(ma)\times D(nb)$. The forms $\lambda$ on $D(A)\times D(B)$ and $\lambda^{\prime\prime}$ on $D(ma)\times D(nb)$ are both split:
@@
[…]
@@ 4.3. Proof of Theorem 2
where $\lambda_{z}=\lambda_{w}$ is a Liouville form on $\mathbb{R}^{2}\backslash\bigl(\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})\bigr)$, with $\Gamma_{1}$ as skeleton. We construct this Liouville form exactly as in Section 3.1, where compactness plays no role. Alternatively, we can do the construction in the proof of Proposition 3.2 on $2,2]^{2}$ in a $\mathbb{Z}^{2}$-periodic way. We then obtain a smooth Liouville form on $0,1]^{2}$ that extends to a smooth $\mathbb{Z}^{2}$-periodic Liouville form on $\mathbb{R}^{2}\backslash\bigl(\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2})\bigr)$ with skeleton $\Gamma_{1}$. $\square$
@@
    Then $\bigl((\Sigma_{N},1),\lambda\bigr)$ is an extendable Liouville polarisation of $N,N[^{4}$ whose skeleton is $\Gamma_{N}$. Notice that $\Sigma_{N}\subset\Sigma_{N+1}$.
@@
    Then $\bigl((\Sigma_{Z},1),\lambda\bigr)$ is a Liouville polarisation of $Z(2)$.
Now $\Sigma_{Z}$ is made of two planes $\{(\pm\frac{1}{2},\frac{1}{2})\}\times\mathbb{R}^{2}$ and infinitely many discs
@@
$\bigl(\{(\Sigma^{\prime}_{1},1),(\Sigma^{\prime}_{2},1)\},\lambda^{\prime}\bigr)$
is a Liouville polarisation of $Z(2)$.
@@
For each $N$ there exists an $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact symplectic morphism $\phi_{N}\colon\widehat{\Sigma}_{N}\to\Sigma^{\prime}$ such that
@@
We also define $\Sigma_{1}^{\prime}(N):=\{(-\frac{1}{2},\frac{1}{2})\}\times[-N,N]^{2}$ and $\Sigma_{2}^{\prime}(N)\subset\Sigma_{2}^{\prime}$ the part of $\Sigma_{2}^{\prime}$ parameterised by $N,N]^{2}\backslash(\mathbb{Z}^{2}+(\frac{1}{2},\frac{1}{2}+\frac{\varepsilon}{2}))$, and notice that they correspond exactly to the intersections of these curves with $Z(2)\cap[-N,N]^{4}$ (see Figure 4.3).
[…]
Figure 4.3. The parametrisation of the curve $\Sigma_{2}^{\prime}$. Each depicted disc $D(\varepsilon)$ has area $\varepsilon$ on the picture, but carries the symplectic form $\omega_{\varepsilon}$ of area $2+\varepsilon$. The black points are the punctures, and the red crosses are the intersection points with $\Sigma_{1}^{\prime}$.
@@
We first construct a symplectic morphism $\varphi_{N}$ by sending a slightly larger disc around the closure of each $\Sigma_{N}^{V}(n,m)$ into $\Sigma_{1}^{\prime}(2N^{2}+1)$, and by sending a slightly larger disc around the closure of each $\Sigma_{N}^{H}(n,m)$ into $\Sigma^{\prime}_{2}(2N^{2}+1)$. The areas of these surfaces are
@@
$\bigl(\Sigma_{1}^{\prime}(2N^{2}+1)\cap\Sigma_{2}^{\prime}(2N^{2}+1)\bigr)=(2(2N^{2}+1))^{2}>16N^{4}=\#(\Sigma_{N}^{V}\cap\Sigma_{N}^{H}).$
We can therefore define $\varphi_{N}$ by its restrictions to $\Sigma_{1}^{V}$ and $\Sigma_{1}^{H}$ as illustrated in Figures 4.4 and 4.5, where $\varphi_{N}$ is explained in blue and $\varphi_{N+1}$ in yellow for $N=1$. These are area preserving embeddings that take the points at the red crosses to points at the red crosses. This defines symplectic morphisms $\varphi_{N}\colon\widehat{\Sigma}_{N}\to\Sigma^{\prime}$, and we can clearly choose $\varphi_{N+1}$ such that $\varphi_{N+1}|_{\widehat{\Sigma}_{N}}=\varphi_{N}$.
The restrictions of these symplectic morphisms to the components of $\widehat{\Sigma}_{N}$, which are discs, are ($\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}}$)-exact. Dealing with the exactness statement is now a matter of applying Lemma 4.2 in a coherent way: This lemma guarantees that $\varphi_{N}$ can be modified to an exact symplectic morphism $\phi_{N}=\varphi_{N}\circ f_{N}$ by precomposition with a Hamiltonian diffeomorphism $f_{N}$ of $\widehat{\Sigma}_{N}$ with compact support. Then $\varphi_{N+1}\circ f_{N}|_{\widehat{\Sigma}_{N}}=\varphi_{N}\circ f_{N}|_{\widehat{\Sigma}_{N}}$, so Remark 4.3 shows that $f_{N+1}$ can be taken to coincide with $f_{N}$ on $\widehat{\Sigma}_{N}$. One now readily checks that the $\phi_{N}$ form a sequence of ($\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}}$)-exact symplectic morphisms such that $\phi_{N+1}|_{\widehat{\Sigma}_{N}}=\phi_{N}$. $\square$
@@
Figure $4.$ The restriction of $\varphi_{1}$ to $\Sigma_{1}^{V}$ and of $\varphi_{2}$ to $\Sigma_{2}^{V}$. Each band has height slightly less than $1$, so that we can construct a larger band of height less than $1$ in the next step $N+1$.
@@
Figure 4.5. The restriction of $\varphi_{1}$ to $\Sigma_{1}^{H}$ and of $\varphi_{2}$ to $\Sigma_{2}^{H}$. These are almost the same embeddings as in the previous figure, except that they avoid the punctures. Note that the area $2+\varepsilon$ of the discs is not even used since for $\varepsilon>0$ small enough the area of $\Sigma_{2}^{\prime}(2N^{2}+1)$ without the discs is already larger than the area of $\Sigma_{N}^{H}$.
@@
[…]
@@
Since the $\phi_{N}$ are $(\alpha_{{\operatorname{st}}}|_{\Sigma_{N}},\alpha_{{\operatorname{st}}}|_{\Sigma^{\prime}})$-exact, $\Psi_{1}$ is $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact. The asserted $(\alpha_{{\operatorname{st}}},\alpha_{{\operatorname{st}}})$-exact symplectic embedding $\Psi\colon{\mathcal{B}}(\Sigma,\lambda)=\mathbb{R}^{4}\setminus\Gamma\to Z(2)$ is now obtained from Lemma 2.21. $\square$